# Towards Optimal Distributed Edge Coloring with Small Palettes

### v1

@@ Abstract
We design a deterministic distributed $\mathcal{O}(\log n)$-round reduction from the $(2\Delta-2)$-edge coloring problem to the much easier $(2\Delta-1)$-edge coloring problem. This is almost optimal, as the $(2\Delta-2)$-edge coloring problem admits an $\Omega(\log_{\Delta}n)$ lower bound. Further, we also obtain an optimal $\mathcal{O}(\log_{\Delta}n)$-round reduction, albeit to the harder maximal independent set (MIS) problem. In our approach we find almost edge-degree-choosable clusters of small diameter. The edges inside these clusters will be colored last by our algorithm. More precisely, we show that if we allow each cluster of uncolored edges to change the color of just two edges in its $1$-hop neighborhood, then it can extend any partial $(2\Delta-2)$-edge coloring. In order to perform these color changes in parallel we need to assign independent edges to each cluster. Therefore we construct an edge set that satisfies two key properties: (1) Each cluster needs to receive exclusive access to two edges in its $1$-hop neighborhood. (2) The set of all such edges needs to be a matching.
Computing a greedy maximal matching satisfies (2), but not necessarily (1). Hence, we modify the maximal matching by letting each cluster send proposals to matching edges in its $2$-hop neighborhood. If the number of proposals that each cluster sends exceeds the number of proposal each matching edge receives we can efficiently assign two matching edges to each cluster. Each cluster can then move its assigned edges into its $1$-hop neighborhood, satisfying (1), while maintaining the matching property (2). This is the only “non-greedy” part of our reduction and it contributes the bulk of the runtime. Finally, each cluster can change the colors of its assigned edges in parallel such that it can extend the coloring to the uncolored edges inside the cluster.
The current state-of-the-art for $(2\Delta-1)$-edge coloring actually comes from an MIS algorithm by [Ghaffari & Grunau, FOCS’24], which runs in $\widetilde{\mathcal{O}}(\log^{5/3}n)$ rounds. With our new reduction, this round complexity now carries over to the $(2\Delta-2)$-edge coloring problem as well. Alternatively, one can also plug in the $(\mathrm{poly}\log\Delta+\mathcal{O}(\log^{\ast}n))$-round $(2\Delta-1)$-edge coloring algorithm from [Balliu, Brandt, Kuhn & Olivetti, PODC’22], which yields an optimal runtime of $\mathcal{O}(\log n)$ rounds for $\Delta\leq\mathrm{poly}\log n$. Previously, the fastest deterministic algorithm using less than $12\Delta-$ 1 colors for general graphs by [Brandt, Maus, Narayanan, Schager & Uitto, SODA’25] ran in $\widetilde{\mathcal{O}}(\log^{3}n)$ rounds. In addition, we also obtain a $\mathcal{O}(\log\log n)$-round randomized reduction of $(2\Delta-2)$-edge coloring to $(2\Delta-1)$-edge coloring. This improves upon the (very recent) best randomized algorithm using less than $12\Delta-$ 1 colors from [Bourreau, Brandt & Nolin, STOC’25] by reducing the round complexity from $\widetilde{\mathcal{O}}(\log^{8/3}\log n)$ down to $\widetilde{\mathcal{O}}(\log^{5/3}\log n)$.
@@ 1 Introduction
In this work we present distributed algorithms for the well-known edge coloring problem. Edge coloring is considered to be one of the four classical problems in the area of local distributed algorithms, together with vertex coloring, maximal independent set (MIS) and maximal matching. A $k$-edge coloring of a graph $G=(V,E)$ is a function $\varphi:E\to\{1,\dots,k\}$ such that no two edges sharing an endpoint receive the same color. The parameter $k$ is called the palette size. Another important parameter for graph coloring problems is the maximum degree $\Delta$. In the centralized setting an edge coloring that uses at least $12\Delta-$ 1 colors can be computed by a trivial greedy algorithm: Simply process all the edges in an arbitrary order and pick any available color at every step. Since any edge is adjacent to most $22\Delta-$ 2 other edges, at least one color will always be available. Similar greedy algorithms also work for the classical $(\Delta+1)$-vertex coloring, MIS and maximal matching problems. Therefore, we will henceforth refer to problems of this type as greedy problems. In contrast, for any palette size $k<2\Delta-1$ there are partial $k$-edge colorings that cannot be extended to the entire graph. In particular, in the online setting, where the input is given as a stream of edges or vertices and the algorithm has to make irrevocable decisions based on partial information on the graph, there is no algorithm that can edge color any graph with less than $12\Delta-$ 1 colors due to a seminal result by Bar-Noy, Motwani and Naor [2]. Therefore this so-called “greedy color threshold” $k=2\Delta-1$ presents a natural and important barrier in the development of distributed algorithms for this problem.
How much more difficult are non-greedy problems than greedy problems?
Going below $12\Delta-$ 1 colors requires completely new techniques, as the algorithm needs to ensure that the yet uncolored edges do not lose too many available colors. This additional difficulty is also demonstrated by the fact that coloring the edges of a graph with less than $12\Delta-$ 1 colors admits a $\Omega(\log_{\Delta}n)$-round deterministic lower bound [9] in the $\mathsf{LOCAL}$ model, while there are deterministic $\mathsf{LOCAL}$ algorithms that compute a $(2\Delta-1)$-edge coloring in $\mathcal{O}(\log^{\ast}n)$ rounds [3, 35], at least in graphs of constant maximum degree. It is known for a long time that every graph can be edge-colored using just $\Delta+1$ colors [43], however until very recently there was no way to compute such a coloring efficiently in the distributed setting.
Model of computation
@@
Our model of computation is the classical message-passing model of distributed computation introduced by Linial in the 1990’s [32, 33]. Today it is known under the apt name $\mathsf{LOCAL}$, coined by Peleg in [40]. In the $\mathsf{LOCAL}$ model the communication network is abstracted as an undirected graph. The vertices are processors with unlimited local computation power and globally unique identifiers. They can communicate with their neighbors in synchronous rounds. The message size is assumed to be unlimited and the complexity of an algorithm is defined as the number of synchronous round until every node outputs (a local part of) the solution. The related $\mathsf{CONGEST}$ model additionally imposes a maximum message size of $\mathcal{O}(\log n)$ bits per message.
Greedy distributed edge coloring
One of the first results for edge coloring in the distributed setting can be derived from Luby’s celebrated MIS algorithm [34]. Via known reductions from edge coloring to vertex coloring to MIS this leads to a randomized algorithm for $(2\Delta-1)$-edge coloring running in $\mathcal{O}(\log n)$ rounds of the $\mathsf{LOCAL}$ model. Further, early works by Panconesi & Srinivasan [36, 39] show that using network decompositions, one can find a $(2\Delta-1)$-edge coloring in $2^{\mathcal{O}(\sqrt{\log n})}$ rounds deterministically. For graphs with constant maximum degree $\Delta$, one can even solve the problem in $\mathcal{O}(\log^{\ast}n)$ rounds [35], matching the famous $\Omega(\log^{\ast}n)$ lower bound by Linial [32, 33]. For general graphs, however, the hunt for the first polylogarithmic-time algorithm that uses just $12\Delta-$ 1 colors was an important and long-standing open question in the field until it was finally positively answered by Fischer, Ghaffari & Kuhn in 2017 [15]. The key technical ingredient of their seminal $\mathcal{O}(\log^{8}\Delta\cdot\log n)$-round algorithm is a deterministic algorithm for the hypergraph maximal matching problem. Since then, the field has picked up the pace with a lot of new algorithms being developed in the last decade, e.g. [1, 19, 20, 21, 25, 29, 30]. The state-of-the-art $(2\Delta-1)$-edge coloring for general graphs now comes out of an MIS algorithm developed in [20]. It runs in $\widetilde{\mathcal{O}}(\log^{5/3}n)$ rounds deterministic and $\widetilde{\mathcal{O}}(\log^{5/3}\log n)$ rounds randomized time.
Breaking the greedy threshold
While most of the algorithms that break the greedy color barrier were only developed in the last ten years, there are two remarkable results before the turn of the millennium. In a seminal paper from 1992, Panconesi & Srinivasan [37, 38] presented the first distributed algorithm for the $\Delta$-vertex coloring algorithm. Their algorithm is deterministic and runs in $2^{\mathcal{O}(\sqrt{\log n})}$ rounds of the $\mathsf{LOCAL}$ model. Since the edge coloring problem on a graph of maximum degree $\Delta$ is equivalent to coloring the vertices of the corresponding line graph of maximum degree $22\Delta-$ 2, this result also yields the first edge-coloring algorithm below the greedy color threshold. On the randomized side, in 1998, Dubhashi, Grable & Panconesi [12] presented an $(1+\varepsilon)\Delta$-edge coloring algorithm running in $\mathcal{O}(\log n)$ rounds, provided that $\Delta>(\log n)^{1+\gamma}$, using the Rödl nibble method. Subsequently, both the number of rounds, as well as the palette size has undergone significant further improvements, which we will discuss in detail in Section 1.3. To put our new results into context we will just briefly mention the relevant state-of-the art algorithms for $(2\Delta-2)$-edge coloring here. In a recent paper studying the hypergraph sinkless orientation problem [8], the authors present a deterministic $(3/2+\varepsilon)\Delta$-edge coloring algorithm running in $\widetilde{\mathcal{O}}(\varepsilon^{-2}\cdot\log^{2}\Delta\cdot\log n)$ rounds. This algorithm is still the fastest for general graphs, even if we allow the palette size to grow as large as $22\Delta-$ 2. On the randomized side, the fastest $(2\Delta-2)$-edge coloring algorithm works by computing a $\Delta$-vertex coloring in $\widetilde{\mathcal{O}}(\log^{8/3}\log n)$ rounds [7].
@@ 1.1 Our contributions
There exists a deterministic distributed algorithm that computes a $(2\Delta-2)$-edge coloring in any $n$-vertex graph with maximum degree $\Delta$ in $\mathcal{O}(\log_{\Delta}n)+T_{\mathrm{MIS}}(\Delta^{2}\cdot n,\mathrm{poly}(\Delta))$ rounds of the $\mathsf{LOCAL}$ model.
By plugging in the state-of-the-art MIS algorithm from [20] we get the following runtime.
@@
There exists a deterministic distributed algorithm that computes a $(2\Delta-2)$-edge coloring in any $n$-vertex graph with maximum degree $\Delta$ in $\widetilde{\mathcal{O}}(\log^{5/3}n)$ rounds of the $\mathsf{LOCAL}$ model.
Secondly, via replacing the expensive MIS by a cheaper ruling set we also get a slightly slower $\mathcal{O}(\log n)$-round reduction to the much easier $(2\Delta-1)$-edge coloring problem.
Theorem 1.3.
There exists a deterministic distributed algorithm that computes a $(2\Delta-2)$-edge coloring in any $n$-vertex graph with maximum degree $\Delta$ in $\mathcal{O}(\log n)+T_{2\Delta-1}(n,\Delta)$ rounds of the $\mathsf{LOCAL}$ model.
Now we can also plug in the $(\mathrm{poly}\log\Delta+\mathcal{O}(\log^{\ast}n))$-round $(2\Delta-1)$-edge coloring algorithm from [1] to get an almost optimal runtime for $\Delta\leq\mathrm{poly}\log n$.
Corollary 1.4.
There exists a deterministic distributed algorithm that computes a $(2\Delta-2)$-edge coloring in any $n$-vertex graph with maximum degree $\Delta$ in $\mathrm{poly}\log\Delta+\mathcal{O}(\log n)$ rounds of the $\mathsf{LOCAL}$ model.
Further, we show that our deterministic result can be modified to also yield a faster randomized reduction. Since our main technical contributions are already fast enough, this only requires us to replace the deterministic subroutines for ruling sets, maximal matchings, etc. by their (sometimes exponentially) faster randomized counterparts.
@@
Again, plugging in the state-of-the-art randomized $(2\Delta-1)$-edge coloring algorithm from [20] yields the following explicit runtime.
@@
There exists a randomized distributed algorithm that computes a $(2\Delta-2)$-edge coloring in any $n$-vertex graph with maximum degree $\Delta$ in $\widetilde{\mathcal{O}}(\log^{5/3}\log n)$ rounds of the $\mathsf{LOCAL}$ model.
@@ 1.2 Our technique in a nutshell
The core observation forming the heart of our reduction is that partial $(2\Delta-2)$-edge colorings can almost always be extended. To give some intuition for this fact, consider the simplest possible case, where the set of uncolored edges in relation to a partial $(2\Delta-2)$-edge coloring $\varphi$ forms a star graph $S\subseteq G$. The only case, in which $\varphi$ cannot be extended to $S$ is if each edge in $S$ has the exact same set of $\Delta-1$ colors appearing among its adjacent edges, which we illustrate in Figure 1. An interesting application of Hall’s marriage theorem shows that this is indeed the only bad case. Next, we will show that this condition generalizes to trees as well. The core observation of this second step is that as long as the leaves in layer $k$ of the tree see at least $\Delta$ distinct colors among their incident edges, we can extend the coloring such that the leaves in layer $k-1$ see at least $\Delta$ distinct colors as well, see Figure 2. Hence, it is sufficient to have one layer of the tree that satisfies this colorful condition.
Lemma 1.7 (Simplified version of Lemma 3.4).
Consider a $(2\Delta-2)$-edge colored graph except for a complete $\Delta$-ary subtree $T\subseteq G$. The coloring can be extended to $T$, if the leaves of $T$ have at least $\Delta$ colors appearing on their incident edges combined.
In the remainder of this paper we will refer to the above condition as the colorful condition. In the full version of this lemma, we generalize the statement to any subtree of a $\Delta$-regular graph, which requires a little bit more care in how one formulates the condition of having enough colors available for the leaves. The main technical difficulty to overcome is that now the leaves of $T$ can all be at different distances from the root.
The reason why this condition is so powerful is because it does not require much intervention in order to fulfil it. For a $\Delta$-regular graph $G$, all we need to do is to change the color of one single edge adjacent to the cluster. Since each leaf in $T$ is already adjacent to $\Delta-1$ colored edges with necessarily distinct colors we only need to obtain one single additional color. However, there are still two big challenges ahead of us, if we want to make this approach work. Firstly, we need to ensure that the chosen edge actually has a useful free color to switch to. Secondly, we need to ensure that all clusters can perform this color switching in parallel without interfering with each other.
To overcome the first obstacle, we construct a subgraph $G^{\prime}\subseteq G$ of maximum degree at most $\Delta-1$ that contains all edges of $G$ except for a spanning tree of each cluster. This subgraph can then be $(2\Delta-3)$-edge coloring using a regular $(2\Delta-1)$-edge coloring algorithm. Hence, every edge in $E(G^{\prime})$ has at least the color $22\Delta-$ 2 available.
To address the second obstacle, we need that the set of edges selected by each cluster is independent, that is, it forms a matching in $G^{\prime}$. Also, for technical reasons we will assign two edges to each cluster instead of just one. Each cluster will only modify the color of one of those two edges, but it needs to ensure that the color of the second edge remains unchanged. As a good first step towards our goal we construct a maximal matching $M$ of $G^{\prime}$, which is still a greedy problem. This is not yet enough though, as the maximality of $M$ only guarantees that each vertex on the boundary of a given cluster sees at least one matching edge in its $2$-hop neighborhood. Hence, there might be a cluster entirely without adjacent matching edges.
@@
Therefore our next step is to modify this matching to meet the above criterion. To solve this problem, we first let the clusters compete for the current edges in $M$. Each cluster $C$ wants to obtain exclusive access to two edges from $M$, which it can then shift along a shortest path towards $C$ until the edge is adjacent to $C$. Hence, each cluster selects a sufficiently large set of vertices on its boundary and lets each such vertex send a proposal to one of its closest matching edges. Note that each edge $e\in M$ may receive proposals from up to $2\Delta$ different clusters. Hence, it suggests itself to model this problem as a hypergraph $H$, where the vertices are the clusters in $G$. Each edge $e\in M$ defines a hyperedge by collecting all clusters that sent a proposals to $e$. What we need to solve now is the non-greedy assignment problem of hyperedges (matching edges) to vertices (clusters). Luckily, this problem has already been studied in the literature, where it is known under the name hypergraph sinkless orientation. An orientation of a hypergraph $H$ defines each hyperedge to be outgoing for exactly one vertex and ingoing for all others. A sinkless orientation on $H$ then requires that every vertex $v$ has at least one outgoing hyperedge, which we can then assign to $v$. As long as each cluster sends out significantly more proposals than any edge in $M$ receives, this problem can be solved in $\mathcal{O}(\log n)$ rounds deterministically and $\mathcal{O}(\log\log n)$ rounds randomized [8]. Finally, each cluster now has two edges, whose colors it can control independently of all other clusters. In Lemma 4.16 we formally show that this is indeed enough to fulfil the colorful condition in Lemma 3.4.
(a) The edges with different colors palettes available are dashed.
Figure 2: Extending the coloring to a tree.

### v2

# Towards Optimal Distributed Edge Coloring with Fewer Colors

@@ Abstract
There is a huge difference in techniques and runtimes of distributed algorithms for problems that can be solved by a sequential greedy algorithm and those that cannot. A prime example of this contrast appears in the edge coloring problem: while $(2\Delta-1)$-edge coloring—where $\Delta$ is the maximum degree—can be solved in $\mathcal{O}(\log^{\ast}(n))$ rounds on constant-degree graphs, the seemingly minor reduction to $(2\Delta-2)$ colors leads to an $\Omega(\log n)$ lower bound [Chang, He, Li, Pettie & Uitto, SODA’18]. Understanding this sharp divide between very local problems and inherently more global ones remains a central open question in distributed computing and it is a core focus of this paper.
As our main contribution we design a deterministic distributed $\mathcal{O}(\log n)$-round reduction from the $(2\Delta-2)$-edge coloring problem to the much easier $(2\Delta-1)$-edge coloring problem. This reduction is optimal, as the $(2\Delta-2)$-edge coloring problem admits an $\Omega(\log n)$ lower bound that even holds on the class of constant-degree graphs, whereas the $12\Delta-$ 1-edge coloring problem can be solved in $\mathcal{O}(\log^{\ast}n)$ rounds. By plugging in the $(2\Delta-1)$-edge coloring algorithms from [Balliu, Brandt, Kuhn & Olivetti, PODC’22] running in $\mathcal{O}(\log^{12}\Delta+\log^{\ast}n)$ rounds, we obtain an optimal runtime of $\mathcal{O}(\log n)$ rounds as long as $\Delta=2^{\mathcal{O}(\log^{1/12}n)}$. Previously, such an optimal algorithm was only known for the class of constant-degree graphs [Brandt, Maus, Narayanan, Schager & Uitto, SODA’25]. Furthermore, on general graphs our reduction improves the runtime from $\widetilde{\mathcal{O}}(\log^{3}n)$ to $\widetilde{\mathcal{O}}(\log^{5/3}n)$.
In addition, we also obtain an optimal $\mathcal{O}(\log\log n)$-round randomized reduction of $(2\Delta-2)$-edge coloring to $(2\Delta-1)$-edge coloring. This leads to a $\widetilde{\mathcal{O}}(\log^{5/3}\log n)$-round $(2\Delta-2$)-edge coloring algorithm, which beats the (very recent) previous state-of-the-art taking $\widetilde{\mathcal{O}}(\log^{8/3}\log n)$ rounds from [Bourreau, Brandt & Nolin, STOC’25].
Lastly, we obtain an $\mathcal{O}(\log_{\Delta}n)$-round reduction from the $(2\Delta-1)$-edge coloring, albeit to the somewhat harder maximal independent set (MIS) problem.
@@ 1 Introduction
There is a huge difference in the techniques and runtimes of distributed algorithms for locally checkable problems that can be solved by a sequential greedy algorithm and those that cannot. For example, on bounded degree graphs, the former problems have distributed complexity $\Theta(\log^{\ast}n)$ [24, 54, 60] while the latter require at least $\Omega(\log n)$ time [17, 19]. Understanding this sharp and potentially wide divide between very locally solvable problems and inherently more global ones remains a central open question in distributed computing.
A prime example of this contrast—explained in detail next and forming the main topic of this paper—appears in the edge coloring problem. A $k$-edge coloring of a graph $G=(V,E)$ is a function $\varphi:E\to\{1,\dots,k\}$ such that no two adjacent edges receive the same color. This problem [60, 25, 28, 31, 4, 16], but also other greedy problems like $(\Delta+1)$-vertex coloring, maximal independent set, and maximal matching [56, 1, 11, 6, 20, 2, 65, 43, 29, 37], has been extensively studied in the field of distributed computing, particularly in the $\mathsf{LOCAL}$ model.
@@
In this model the communication network is abstracted as an undirected graph $G=(V,E)$. Each vertex hosts a processor that may communicate with its neighbors in synchronous rounds. In one round, each processor may send unbounded-size messages to all its neighbors in $G$ and perform unbounded local computations. The goal is to minimize the number of synchronous rounds until every node outputs (a local part of) the solution. The related $\mathsf{CONGEST}$ model additionally imposes a maximum message size of $\mathcal{O}(\log n)$ bits per message.
In order to make progress in many parts of the graph in parallel, many distributed algorithms crucially rely on the property that color choices made by other vertices do not affect the solvability on the remaining graph. While this property is satisfied for $k=2\Delta-1$ (hence we call this the greedy threshold), for any $k<2\Delta-1$ there are partial $k$-edge colorings that cannot be extended to the entire graph [5]. On the other hand of the spectrum, it is known that every graph admits a $(\Delta+1)$-edge coloring by a celebrated result of Vizing [69].
The intuition that coloring with fewer colors is significantly more challenging—particularly in the distributed setting—is evidenced by strong lower bounds: Any deterministic edge coloring algorithm using less than $12\Delta-$ 1 colors requires at least $\Omega(\log_{\Delta}n)$ rounds, and even randomized algorithms face an $\Omega(\log_{\Delta}\log n)$-round lower bound [17]. On trees, where these lower bounds are established, there is even a deterministic algorithm that solves the greedy version of the problem in strongly sublogarithmic time, specifically in $\mathcal{O}(\log^{12/13}n)$ rounds [16]. On bounded degree graphs, the gap is even larger as the greedy version has deterministic and randomized complexity $\Theta(\log^{\ast}n)$, yet the lower bounds for algorithms that use fewer colors—$\Omega(\log n)$ deterministic and $\Omega(\log\log n)$ randomized—still apply.
Despite these challenges, there has been extensive work on designing increasingly faster deterministic and randomized algorithms for edge coloring with fewer than $12\Delta-$ 1 colors—both in the $\mathsf{LOCAL}$ model [12, 13, 15, 17, 18, 44, 49], the more restricted $\mathsf{CONGEST}$ model [44, 47], and in other models such as streaming [5, 23, 66]. The fastest known deterministic distributed algorithm that operates below the greedy threshold runs in $\widetilde{\mathcal{O}}(\log^{3}n)$ rounds [^1] We use the notation $\widetilde{\mathcal{O}}(f(n))$ to hide factors polylogarithmic in $f(n)$. [15], while the fastest randomized algorithm, derived from a recent more general result on $\Delta$-vertex coloring, achieves a runtime of $\widetilde{\mathcal{O}}(\log^{8/3}\log n)$ rounds [14]. See the related work section (Section 1.3) for further results on edge coloring with fewer colors, and refer to Table 1 for an overview of existing algorithms, the number of colors they use, and their runtime.
This work aims to better understand the size and nature of the complexity gap between greedy-solvable problems and those below the greedy threshold. Understanding this latter class of problems is also central to the poorly understood regime in the complexity landscape of local graph problems, see, e.g., [30, 22, 58].
@@ 1.1 Our contributions
There is a deterministic $\mathsf{LOCAL}$ algorithm computing a $(2\Delta-2)$-edge coloring in any $n$-vertex graph with maximum degree $\Delta$ in $\mathcal{O}(\log n)+T_{2\Delta-1}(n,\Delta-1)$ rounds.
Secondly, allowing the use of maximal independent set (MIS) as a subroutine improves the runtime of our reduction at the cost of reducing to a harder problem.
Theorem 1.2.
There is a deterministic $\mathsf{LOCAL}$ algorithm computing a $(2\Delta-2)$-edge coloring in any $n$-vertex graph with maximum degree $\Delta$ in $\mathcal{O}(\log_{\Delta}n)+T_{\mathrm{MIS}}(\Delta^{2}\cdot n,\mathrm{poly}(\Delta))$ rounds.
While the latter result might sound like a strict improvement over the former—especially given the fact that the runtimes of the current state-of-the-art algorithms for $(2\Delta-1)$-edge coloring and MIS on general graphs coincide—we argue that the former version is, in fact, the stronger result. A classic reduction from Luby [56] shows that $(2\Delta-1)$-edge coloring is no harder than MIS. Consequently, the $\widetilde{\mathcal{O}}(\log^{5/3}n)$ deterministic algorithm for MIS [37] also solves the $(2\Delta-1)$-edge coloring problem. On the other hand, there is a $\mathcal{O}(\log^{12}\Delta+\log^{\ast}n)$-round $(2\Delta-1)$-edge coloring algorithm from [4], which provably cannot be achieved by any MIS algorithm due to the $\Omega(\min(\Delta,\log n/\log\log n))$ lower bound established in [3]. We obtain the following corollary.
@@
There are deterministic $\mathsf{LOCAL}$ algorithms computing a $(2\Delta-2)$-edge coloring in any $n$-vertex graph with maximum degree $\Delta$ within
  1. 1.
    either $\mathcal{O}(\log^{12}\Delta+\log n)$ rounds,
  2. 2.
    or $\widetilde{\mathcal{O}}(\log^{5/3}n)$ rounds.
For the class of graphs with maximum degree $\Delta$ as large as $2^{\mathcal{O}(\log^{1/12}n)}$, the first point of this corollary yields an optimal runtime of $\mathcal{O}(\log n)$ rounds. Previously, an optimal runtime was only known for the class of bounded-degree graphs [15]. The second point significantly improves the upper bound on the complexity of $(2\Delta-2)$-edge coloring on general graphs, which previously stood at $\widetilde{\mathcal{O}}(\log^{3}n)$ rounds [15]. Additionally, we present an exponentially faster randomized reduction to the $(2\Delta-1)$-edge coloring problem.
@@
Theorem 1.4 yields the following corollary by plugging in the algorithms of [4] and [37].
@@
There are randomized distributed algorithms computing a $(2\Delta-2)$-edge coloring in any $n$-vertex graph with maximum degree $\Delta$ within
  1. 1.
    either $\mathcal{O}(\log^{12}\Delta+\log\log n)$ rounds,
  2. 2.
    or $\widetilde{\mathcal{O}}(\log^{5/3}\log n)$ rounds of the $\mathsf{LOCAL}$ model.
For the class of graphs with maximum degree $\Delta$ as large as $2^{\mathcal{O}(\log^{1/12}\log n)}$, the first point of this corollary yields an optimal runtime of $\mathcal{O}(\log\log n)$ rounds. The second point beats the (very recent) previous state-of-the-art for randomized $(2\Delta-2)$-edge coloring on general graphs, which previously stood at $\widetilde{\mathcal{O}}(\log^{8/3}\log n)$ rounds [14].
@@ 1.2 Our technique in a nutshell
To illustrate the difficulty of the $(2\Delta-2)$-edge coloring problem, let us first try to design an $\mathcal{O}(1)$-round reduction to the MIS and the $(2\Delta-1)$-edge coloring problem; recall that such a reduction provably cannot exist. One approach is to use the MIS algorithm to compute a clustering of the graph, e.g., by computing an MIS of the power graph $G^{2}$ and letting each vertex join the cluster of the closest node in the MIS; see Figure 1 for an example. As a result, we can partition the edge set of $G$ into edges $E_{\mathrm{intra}}$ (black) within the clusters and inter-cluster edges $E_{\mathrm{inter}}$ (turquoise) that go between the different clusters. We could simply color the edges in $E_{\mathrm{intra}}$ independently and in parallel for each cluster, as the edges of different clusters are non-adjacent. Due to the constant diameter of each cluster, this can be done in $O(1)$ rounds and even using the existentially optimal number of colors, i.e., $\Delta+1$ colors. However, now the problem of coloring $E_{\mathrm{intra}}$ is no easier than the original $(2\Delta-2)$-edge coloring problem as it essentially corresponds to coloring a general graph but with some colors—the ones used by adjacent edges in $E_{\mathrm{intra}}$—being forbidden for certain edges. In fact, it might not even be possible to complete the coloring at all.
Figure 1: A matching of the intercluster edges that assigns two edges to each clusters. The matching edges are marked red. Arrows point towards the cluster the edge belongs to.
Instead, in our algorithm, we first color the edges between the clusters and then try to complete it within the clusters. First, observe that the graph induced by $E_{\mathrm{inter}}$ has maximum degree $\Delta^{\prime}=\Delta-1$, as each node has at least one adjacent edge within a cluster. Thus, we can color $E_{\mathrm{inter}}$ with one instance of greedy-type edge coloring, using only $32\Delta^{\prime}-1=2\Delta-$ 3 colors. The main challenge that we address in this paper is how to ensure that the coloring can be extended to the edges within the clusters.
For the rest of the this technical overview we assume that the graph is $\Delta$-regular with a sufficiently high girth such that each cluster looks like a tree; our full proof does not require this assumption. In addition, let us assume for now that each cluster is a perfect $\Delta$-ary tree, that is, all leaves are at the same level. Then, a simplified version of our main technical contribution gives a useful condition under which we can complete the coloring.
Lemma 1.6 (Simplified version of Lemma 3.4).
Consider a $(2\Delta-2)$-edge colored graph except for a perfect $\Delta$-ary subtree $T\subseteq G$. If the number of distinct colors appearing on the incident edges of any leaf in $T$ combined is at least $\Delta$, then the coloring can be extended to $T$.
For the remainder of this paper we will refer to the above condition as the colorful condition. To build some intuition for it, consider the simplest possible case, where the set of uncolored edges forms a star graph $S\subseteq G$. The only case, in which $\varphi$ cannot be extended to $S$ is if each edge in $S$ has the exact same set of $\Delta-1$ colors appearing among its adjacent edges, which we illustrate in Figure 2. By setting up a bipartite graph with available colors on one side and edges on the opposite side, an application of Hall’s marriage theorem shows that this is indeed the only bad case. In the proof of Lemma 3.4, we generalize this condition from uncolored stars to uncolored trees.
@@
Next, we turn our attention to how the partial edge coloring of $E_{\mathrm{inter}}$ can be modified in order to satisfy the colorful condition for every cluster of uncolored edges. The core observation is that each leaf in the cluster already has $\Delta-1$ incident edges in $E_{\mathrm{inter}}$, whose colors must all be distinct. Therefore, if we can manage to keep the set of colors fixed for one leaf and modify at most one color at another leaf, the colorful condition is already satisfied. Notably, our partial coloring of $E_{\mathrm{inter}}$ only uses $32\Delta-$ 3 of the $22\Delta-$ 2 available colors. As a result, every edge already has at least one free color it can switch to if needed.
The main difficulty now is to construct a set of edges such that each cluster can perform these changes in parallel and without causing any conflicts. Hence, our desired set of edges needs to satisfy two key properties: (1) Each cluster needs exclusive access to two edges in its $1$-hop neighborhood. (2) The set of all such edges needs to be a matching. Computing a greedy maximal matching satisfies (2), but not necessarily (1). Therefore, we modify a maximal matching $M\subseteq E_{\mathrm{inter}}$ by letting each cluster send proposals to matching edges in its $2$-hop neighborhood.
(a) The intracluster edges $E_{\mathrm{intra}}$ are dashed. We first compute a maximal matching $M\subseteq E_{\mathrm{inter}}$, which is highlighted red.
(b) Next, we assign two edges in $M$ to each cluster. This is the only “non-greedy” part of our reduction.
(c) Finally, we show how each cluster can move those edges into its neighborhood while maintaining the matching property.
Figure 3: In order to complete the coloring on the edges within the clusters, our algorithm ensures that each cluster can independently control the colors of two edges adjacent to the cluster. For illustration purposes this example has many cycles that are dealt with separately in our algorithm.
@@
To solve this non-greedy problem of assigning matching edges to clusters, we set up a hypergraph $H$. Each cluster in $G$ corresponds to a vertex in $H$ and each edge $e\in M$ defines a hyperedge containing all clusters that sent at least one proposal to $e$. Note that a hypergraph is in fact necessary here, since each edge in $M$ may receive proposals from up to $2\Delta$ different clusters. Now our assignment problem can then be viewed as a hypergraph orientation problem, where each hyperedge is said to be outgoing for exactly one of its vertices and ingoing for all others. In this setting, the problem has already been studied in the literature under the name hypergraph sinkless orientation. As long as each cluster sends out sufficiently more proposals than any edge in $M$ receives, it can be solved in $\mathcal{O}(\log_{\Delta}n)$ rounds deterministically and $\mathcal{O}(\log_{\Delta}\log n)$ rounds randomized [15]. This is the only “non-greedy” part of our reduction and it dominates the runtime. Every other step runs in $\mathcal{O}(\log^{\ast}n)$ rounds for constant-degree graphs, which shows that this is an essential part of our reduction in light of the $\Omega(\log_{\Delta}n)$ lower bound for edge-coloring with less than $12\Delta-$ 1 colors.
Now as a final technical remark, while each cluster may own two assigned edges, these may not be directly adjacent to the cluster and controlling their color would still be useless in terms of obtaining solvability of the cluster. To solve this issue we will show that cluster $C$ can move its assigned edges into its $1$-hop neighborhood along a shortest path to $C$, satisfying (1), while maintaining the matching property (2). Finally, the colorful condition is satisfied for every cluster and we can complete the $(2\Delta-2)$-edge coloring for each cluster in parallel.